# A parallel domain decomposition method for solving elliptic equations on manifolds

### v1

@@ 3.2 Numerical Algorithm
$\rho_{i}=\frac{\sigma_{i}}{\sum_{j=1}^{m}\sigma_{j}}.$ (15)
@@
is a diffeomorphism which is the transition of coordinates on the overlapping between $M_{i}$ and $M_{j}$. We refer to [10, § 3.2] for a detailed explanation of the transition maps. (Particularly, Fig. 1 therein provides an illustration.) We still wish to point out that $\phi_{j}^{-1}\circ\phi_{i}$ preserves neither nodes nor grid necessarily. In other words, $\phi_{j}^{-1}\circ\phi_{i}$ may neither map a node in $D_{ij}$ to a node in $D_{ji}$, nor map the grid over $D_{ij}$ to the one over $D_{ji}$. As emphasized in [10, § 3.2], this incompatibility among the grids over different $D_{i}$ actually shows the high flexibility of our approach.
@@
The computation in (17) needs more explanation. If $\phi_{i}(\xi)\notin M_{j}$, then $\rho_{j}(\phi_{i}(\xi))=0$ and the summand $\rho_{j}(\phi_{i}(\xi))\cdot u_{h,j}^{n-1}(\phi_{j}^{-1}\circ\phi_{i}(\xi))$ is actually not needed. On the contrary, $\phi_{i}(\xi)\in M_{j}$,
@@ 4 Manifolds with Boundary
such that $V$ is a neighborhood of $\partial M$ in $M$. Such a $V$ is called a collar neighborhood of $\partial M$. One may even arrange $\psi(\{\delta\}\times\partial M)=\partial M$ or $\psi(\{\eta\}\times\partial M)=\partial M$ as one wishes. Let’s assume $\psi(\{\eta\}\times\partial M)=\partial M$. (See Fig. 4 for an illustration, the disk therein is a manifold with boundary, the part with segments is a collar neighborhood $\psi((\delta,\eta]\times\partial M)$, each segment is parameterized by $(\delta,\eta]$.)
Figure 4: An illustration of collar neighborhood.
@@ 5 Numerical Experiments
We perform several numerical tests of Algorithm 2. The proposed method certainly applies to problems in all dimensions. However, the sizes of the linear systems derived from subdomains will increase exponentially with respect to the dimension. This difficulty is so called “the curse of dimensionality”. Due to the constraint of computing resources, just as in [10], we only deal with problems with dimension no more than $4$. Nevertheless, the numerical examples in [10] are all on manifolds without boundary. We now deal with manifolds with boundary too.
Our examples of manifolds are $\mathbb{CP}^{2}$, $S^{2}\times S^{2}$, $B^{4}$ and $B^{2}\times S^{2}$. All of them are of dimension $4$, the former two are without boundary, and the latter two are with boundary. It’s also worth noting that, if a manifold can be decomposed into two subdomains, say the $d$-dimensional unit sphere $S^{d}$, the proposed method coincides essentially with the Schwarz Alternating Method and the method in [10]. As a result, we do not perform test on $S^{4}$ now since such an experiment has been done in [10].
@@ 5.1 A Problem on $\mathbb{CP}^{2}$
In this example, the definition of $\mathbb{CP}^{2}$, the Riemannian metric, the concrete problem of (3), the decomposition of $\mathbb{CP}^{2}$, and the transition maps are the same as those in [10, § 5.2]. We refer to [10] for a thorough and detailed description. Here we only briefly overview the decomposition.
Now $\mathbb{CP}^{2}$ is decomposed into $3$ subdomains $M_{j}$ for $0\leq j\leq 2$. (Note that, following the convention of algebraic geometry, the index $j$ is chosen to start from 00.) In the following, $z_{j}=x_{j}+\sqrt{-1}y_{j}\in\mathbb{C}$, $x_{j}\in\mathbb{R}$, and $y_{j}\in\mathbb{R}$. We shall identify the complex number $z_{j}$ with the $2$-dimensional real vector $(x_{j},y_{j})$. Let $D_{j}=[-r,r]^{4}\subset\mathbb{R}^{4}\simeq\mathbb{C}^{2}$. For $0\leq j\leq 2$, we have the following diffeomorphisms
$\phi_{0}:\ D_{0}$$\rightarrow$ $M_{0}\subset\mathbb{CP}^{2}$
$(z_{1},z_{2})$ $\mapsto$ $[1,z_{1},z_{2}],$
$\phi_{1}:\ D_{1}$$\rightarrow$ $M_{1}\subset\mathbb{CP}^{2}$
$(z_{0},z_{2})$ $\mapsto$ $[z_{0},1,z_{2}],$
and
$\phi_{2}:\ D_{2}$$\rightarrow$ $M_{2}\subset\mathbb{CP}^{2}$
$(z_{0},z_{1})$ $\mapsto$ $[z_{0},z_{1},1],$
where $[w_{0},w_{1},w_{2}]$ are called the homogeneous coordinates of $\mathbb{CP}^{2}$. (For the definition of homogeneous coordinates, see [10].) To guarantee $\mathbb{CP}^{2}=\bigcup_{j=0}^{2}(M_{j}\setminus\partial M_{j})$, we have to let $r>1$. The larger $r$ is, the more overlapping there will be.
@@
$\sigma_{j}\circ\phi_{j}(x)=\begin{cases}0,&\max\{|x_{j}|\mid 1\leq j\leq 4\}>r^{\prime};\\ \prod_{j=1}^{4}(1-(\tfrac{x_{j}}{r^{\prime}})^{2}),&\text{otherwise}.\end{cases}$ (18)
It’s easy to check that $\mathrm{supp}\sigma_{j}=\phi_{j}([-r^{\prime},r^{\prime}]^{4})\subset M_{j}\setminus\partial M_{j}$, $\sigma_{j}$ is Lipschitz continuous, and $\sum_{j=0}^{2}\sigma_{j}>0$. Hence the resulted family of $\rho_{j}$ satisfy Assumption 2.1.
For the discretization of $D_{j}=[-r,r]^{4}$, we divide each coordinate interval $[-r,r]$ into $N$ equal parts. The scale of the grid is thus $h=2r/N$. There are $(N+1)^{4}$ nodes on $D_{j}$, most rows of the stiffness matrix have $813^{4}=$ 81 nonzero entries. We keep $N\leq 80$ due to the memory limitation of the hardware.
@@
We see that the error $I_{h}u-u_{h}^{\infty}$ decays in the optimal order when $h$ decreases. Furthermore, $n_{0}$ decreases when $r$ becomes larger, i.e. $u_{h}^{n}$ reaches its limit $u_{h}^{\infty}$ fast provides that the overlapping between subdomains is large.
5.2 A Problem on $S^{2}\times S^{2}$
Let $M=S^{2}\times S^{2}$, where $S^{2}$ is the unit sphere in $\mathbb{R}^{3}$, i.e.
$S^{2}=\left\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}\middle|\sum_{i=1}^{3}y_{i}^{2}=1\right\}.$
In this example, the Riemannian metric, the concrete problem of (3), the decomposition of $M$, and the transition maps are the same as those in [10, § 5.3]. We refer to [10] for a thorough and detailed description. Here we only briefly overview the decomposition.
We decompose $S^{2}\times S^{2}$ into $4$ subdomains. In the following, let $x=(x_{1},x_{2})$ and $x^{\prime}=(x^{\prime}_{1},x^{\prime}_{2})$ denote the coordinates of $\mathbb{R}^{2}$. Let $\|x\|=\sqrt{\sum_{i=1}^{2}x_{i}^{2}}$ and $\|x^{\prime}\|=\sqrt{\sum_{i=1}^{2}{x^{\prime}}_{i}^{2}}$. For $1\leq i\leq 4$, let
$D_{i}=[-r,r]^{4}=\{(x,x^{\prime})\mid x\in[-r,r]^{2},x^{\prime}\in[-r,r]^{2}\}.$
For $1\leq i\leq 4$, we have the following diffeomorphisms
$\phi_{1}:\ D_{1}$$\rightarrow$ $M_{1}\subset S^{2}\times S^{2}\subset\mathbb{R}^{3}\times\mathbb{R}^{3}$
$(x,x^{\prime})$ $\mapsto$ $\left(\frac{2x}{1+\|x\|^{2}},\frac{1-\|x\|^{2}}{1+\|x\|^{2}},\frac{2x^{\prime}}{1+\|x^{\prime}\|^{2}},\frac{1-\|x^{\prime}\|^{2}}{1+\|x^{\prime}\|^{2}}\right),$
$\phi_{2}:\ D_{2}$$\rightarrow$ $M_{2}\subset S^{2}\times S^{2}\subset\mathbb{R}^{3}\times\mathbb{R}^{3}$
$(x,x^{\prime})$ $\mapsto$ $\left(\frac{2x}{1+\|x\|^{2}},\frac{1-\|x\|^{2}}{1+\|x\|^{2}},\frac{2x^{\prime}}{1+\|x^{\prime}\|^{2}},\frac{-1+\|x^{\prime}\|^{2}}{1+\|x^{\prime}\|^{2}}\right),$
$\phi_{3}:\ D_{3}$$\rightarrow$ $M_{3}\subset S^{2}\times S^{2}\subset\mathbb{R}^{3}\times\mathbb{R}^{3}$
$(x,x^{\prime})$ $\mapsto$ $\left(\frac{2x}{1+\|x\|^{2}},\frac{-1+\|x\|^{2}}{1+\|x\|^{2}},\frac{2x^{\prime}}{1+\|x^{\prime}\|^{2}},\frac{1-\|x^{\prime}\|^{2}}{1+\|x^{\prime}\|^{2}}\right),$
and
$\phi_{4}:\ D_{4}$$\rightarrow$ $M_{4}\subset S^{2}\times S^{2}\subset\mathbb{R}^{3}\times\mathbb{R}^{3}$
$(x,x^{\prime})$ $\mapsto$ $\left(\frac{2x}{1+\|x\|^{2}},\frac{-1+\|x\|^{2}}{1+\|x\|^{2}},\frac{2x^{\prime}}{1+\|x^{\prime}\|^{2}},\frac{-1+\|x^{\prime}\|^{2}}{1+\|x^{\prime}\|^{2}}\right).$
To guarantee $S^{2}\times S^{2}=\bigcup_{i=1}^{4}(M_{i}\setminus\partial M_{i})$, we have to let $r>1$. The larger $r$ is, the more overlapping there will be.
To obtain a partition of unity $\{\rho_{i}\mid 1\leq i\leq 4\}$, it suffices to define the $\sigma_{j}$ in (15). We define $\sigma_{j}$ as (18) again. The grids on $D_{i}$ are the same as those for $\mathbb{CP}^{2}$. We choose the initial guess as 00. The numerical results are as follows in Tables 3 and 4.
$0.24$ $0.0192$ $0.0503$ $0.1444$ $0.2093$ $21$
$0.12$ $0.0043$ $2.2$ $0.0126$ $2.0$ $0.0410$ $1.8$ $0.0577$ $1.9$ $21$
$0.06$ $0.0011$ $2.0$ $0.0031$ $2.0$ $0.0115$ $1.8$ $0.0154$ $1.9$ $21$
$0.03$ $42.8668e-$ 4 $1.9$ $47.6697e-$ 4 $2.0$ $0.0041$ $1.5$ $0.0050$ $1.6$ $21$
Table 3: Numerical result on $S^{2}\times S^{2}$ for $r=1.2$.
$0.4$ $0.1381$ $0.6327$ $0.9428$ $0.9869$ $9$
$0.2$ $0.0228$ $2.6$ $0.1331$ $2.2$ $0.3062$ $1.6$ $0.2908$ $1.8$ $8$
$0.1$ $0.0085$ $1.4$ $0.0363$ $1.9$ $0.0996$ $1.6$ $0.0857$ $1.8$ $8$
$0.05$ $0.0015$ $2.5$ $0.0084$ $2.1$ $0.0329$ $1.6$ $0.0260$ $1.7$ $9$
Table 4: Numerical result on $S^{2}\times S^{2}$ for $r=2$.
@@ 5.3 A Problem on $B^{4}$
be a collar mapping, where $10<\delta<$ 1. (Again, see Fig. 4 for an illustration.) Let $M_{1}=[-s,s]^{4}$, we have
@@
In the above, we have $0.50<\delta<s<$ 0.5 and $r>1$. The larger $s-\delta$ and $r$ are, the more overlapping there will be.
@@
We choose $N_{2}$ as $10$, $20$, $40$, and $80$. For the initial guess, we set the degrees of freedom off $\partial M$ as 00. The numerical results are as follows in Tables 5 and 6.
@@ 5.4 A Problem on $B^{2}\times S^{2}$
Note that $D_{1,1}=D_{1,2}=[-s,s]^{2}\times[-r,r]^{2}$ and other $D_{i,i^{\prime}}$ are $[\delta,1]\times[-r,r]^{3}$. Here $0<\delta<s<\frac{1}{\sqrt{2}}$ and $r>1$. The larger $s-\delta$ and $r$ are, the more overlapping there will be.
@@
Now we choose $b=1$ in (3). For the discretization on $D_{i,i^{\prime}}$, we divide each coordinate interval $[-s,s]$ and $[\delta,1]$ into $N_{1}$ equal parts, divide each coordinate interval $[-r,r]$ into $N_{2}$ equal parts. Set $N_{1}=0.4N_{2}$. We choose $N_{2}$ as $10$, $20$, $40$, and $80$. For the initial guess, we set the degrees of freedom off $\partial M$ as 00. The numerical results are as follows in Tables 7 and 8.
@@ Acknowledgements
We thank Jun Li, Yuanming Xiao and Yiyan Xu for various discussions. The first author was partially supported by NSFC 11871272. The second author was partially supported by NSFC 12071227 and 12371369, and the National Key Research and Development Program of China 2020YFA0713803. The third author was partially supported by NSFC 12171349 and 12271389.

### v2

@@ 3.2 Numerical Algorithm
(17) $\rho_{i}=\frac{\sigma_{i}}{\sum_{j=1}^{m}\sigma_{j}}.$
@@
is a diffeomorphism which is the transition of coordinates on the overlap between $M_{i}$ and $M_{j}$. We refer to [10, § 3.2] for a detailed explanation of the transition maps. (Particularly, Fig. 1 in [10] provides an illustration.) We still wish to point out that $\phi_{j}^{-1}\circ\phi_{i}$ preserves neither nodes nor grid necessarily. In other words, $\phi_{j}^{-1}\circ\phi_{i}$ may neither map a node in $D_{ij}$ to a node in $D_{ji}$, nor map the grid over $D_{ij}$ to the one over $D_{ji}$. As emphasized in [10, § 3.2], this incompatibility among the grids over different $D_{i}$ actually shows the high flexibility of our approach.
@@
The computation in (19) needs more explanation. If $\phi_{i}(\xi)\notin M_{j}$, then $\rho_{j}(\phi_{i}(\xi))=0$ and the summand $\rho_{j}(\phi_{i}(\xi))\cdot u_{h,j}^{n-1}(\phi_{j}^{-1}\circ\phi_{i}(\xi))$ is actually not needed. Otherwise, $\phi_{i}(\xi)\in M_{j}$ and
@@ 4 Manifolds with Boundary
such that $V$ is a neighborhood of $\partial M$ in $M$. Such a $V$ is called a collar neighborhood of $\partial M$. One may even arrange $\psi(\{\delta\}\times\partial M)=\partial M$ or $\psi(\{\eta\}\times\partial M)=\partial M$ as one wishes. Let’s assume $\psi(\{\eta\}\times\partial M)=\partial M$.
@@
To obtain a partition of unity, by (17), it suffices to get certain $\sigma_{i}$. We can construct a $\sigma_{m_{1}+j}$ on $M_{m_{1}+j}$ as follows. Firstly, choose a nonnegative function $\sigma^{\prime}_{j}$ on $M^{\prime}_{j}$ such that $\mathrm{supp}\sigma^{\prime}_{j}\subset M^{\prime}_{j}\setminus\partial M^{\prime}_{j}$. Note that $\sigma^{\prime}_{j}$ is defined on a subdomian $M^{\prime}_{j}$ of $\partial M$, and hence it is a function of $(d-1)$ variables which can be expressed locally by $(d-1)$ coordinates. Choose a nonnegative $1$-variable function $\sigma^{\prime\prime}$ on $[\delta,\eta]$ such that $\mathrm{supp}\sigma^{\prime\prime}=[\delta^{\prime\prime},\eta]$, where $\delta<\delta^{\prime\prime}<\eta$. Then $\sigma^{\prime\prime}\times\sigma^{\prime}_{j}$ is a function on $[\delta,\eta]\times M^{\prime}_{j}\subset[\delta,\eta]\times\partial M$. Define $\sigma_{m_{1}+j}=(\sigma^{\prime\prime}\times\sigma^{\prime}_{j})\circ\psi^{-1}$. We see $\sigma_{m_{1}+j}$ is a nonnegative function on $M_{m_{1}+j}$ with $\mathrm{supp}\sigma_{m_{1}+j}\subset M_{m_{1}+j}\setminus\overline{\gamma_{m_{1}+j}}$.
@@ 5 Numerical Experiments
We perform several numerical tests of Algorithm 2. The proposed method certainly applies to problems in all dimensions. However, the sizes of the linear systems derived from subdomains will increase exponentially with respect to the dimension. This difficulty is so called “the curse of dimensionality”. Due to the constraint of computing resources, just as in [10], we only deal with problems with dimension no more than $4$. Nevertheless, the numerical examples in [10] are all on manifolds without boundary. We now deal with manifolds with boundary too. Our examples of manifolds are $\mathbb{CP}^{2}$, $B^{4}$ and $B^{2}\times S^{2}$. All of them are of dimension $4$, the $\mathbb{CP}^{2}$ is without boundary, and the latter two are with boundary.
@@ 5.1 A Problem on $\mathbb{CP}^{2}$
In this example, the definition of $\mathbb{CP}^{2}$, the Riemannian metric, the concrete problem of (4), the decomposition of $\mathbb{CP}^{2}$, and the transition maps are the same as those in [10, § 5.2]. We refer to [10] for a thorough and detailed description.
Particularly, $\mathbb{CP}^{2}$ is decomposed into $3$ subdomains $M_{j}$ for $0\leq j\leq 2$, which is provided by diffeomorphisms $\phi_{j}:D_{j}\rightarrow M_{j}$. We have $D_{j}=[-r,r]^{4}\subset\mathbb{C}^{2}=\mathbb{R}^{4}$,
$\phi_{0}(z_{1},z_{2})=[1,z_{1},z_{2}],\quad\phi_{1}(z_{0},z_{2})=[z_{0},1,z_{2}],\quad\phi_{2}(z_{0},z_{1})=[z_{0},z_{1},1].$
Here each $z_{j}$ is a complex number which also represents a vector in $\mathbb{R}^{2}$, and $[a,b,c]$ represents the homogenous coordinates of $\mathbb{CP}^{2}$. To guarantee $\mathbb{CP}^{2}=\bigcup_{j=0}^{2}(M_{j}\setminus\partial M_{j})$, we have to let $r>1$. The larger $r$ is, the more overlap there will be.
@@
(20) $\sigma_{j}\circ\phi_{j}(x)=\begin{cases}0,&\max\{|x_{j}|\mid 1\leq j\leq 4\}>r^{\prime};\\ \prod_{j=1}^{4}(1-(\tfrac{x_{j}}{r^{\prime}})^{2}),&\text{otherwise}.\end{cases}$
Indeed, we can firstly define a function on $[-r,r]$ as
$w(a)=\begin{cases}0,&|a|>r^{\prime};\\ 1-(\tfrac{a}{r^{\prime}})^{2},&\text{otherwise}.\end{cases}$
Then $\sigma_{j}\circ\phi_{j}(x)=\prod_{j=1}^{4}w(x_{j})$. We see that $\mathrm{supp}\sigma_{j}=\phi_{j}([-r^{\prime},r^{\prime}]^{4})\subset M_{j}\setminus\partial M_{j}$, $\sigma_{j}$ is Lipschitz continuous, and $\sum_{j=0}^{2}\sigma_{j}>0$. Hence the resulted family of $\rho_{j}$ satisfy Assumption 2.1.
For the discretization of $D_{j}=[-r,r]^{4}$, we divide each coordinate interval $[-r,r]$ into $N$ equal parts. The scale of the grid is thus $h=2r/N$. We keep $N\leq 80$ due to the memory limitation of the hardware.
@@
We see that the error $I_{h}u-u_{h}^{\infty}$ decays in the optimal order when $h$ decreases.
@@ 5.2 A Problem on $B^{4}$
be a collar mapping, where $10<\delta<$ 1. Let $M_{1}=[-s,s]^{4}$, we have
@@
In the above, we have $0.50<\delta<s<$ 0.5 and $r>1$. The larger $s-\delta$ and $r$ are, the more overlap there will be.
@@
We choose $N_{2}$ as $10$, $20$, $40$, and $80$. For the initial guess, we set the degrees of freedom off $\partial M$ as 00. The numerical results are as follows in Tables 3 and 4.
@@ 5.3 A Problem on $B^{2}\times S^{2}$
Note that $D_{1,1}=D_{1,2}=[-s,s]^{2}\times[-r,r]^{2}$ and other $D_{i,i^{\prime}}$ are $[\delta,1]\times[-r,r]^{3}$. Here $0<\delta<s<\frac{1}{\sqrt{2}}$ and $r>1$. The larger $s-\delta$ and $r$ are, the more overlap there will be.
@@
Now we choose $b=1$ in (4). For the discretization on $D_{i,i^{\prime}}$, we divide each coordinate interval $[-s,s]$ and $[\delta,1]$ into $N_{1}$ equal parts, divide each coordinate interval $[-r,r]$ into $N_{2}$ equal parts. Set $N_{1}=0.4N_{2}$. We choose $N_{2}$ as $10$, $20$, $40$, and $80$. For the initial guess, we set the degrees of freedom off $\partial M$ as 00. The numerical results are as follows in Tables 5 and 6.
@@ Acknowledgements
We thank Jun Li, Yuanming Xiao and Yiyan Xu for various discussions. We thank the anonymous reviewers for many constructive comments and suggestions which led to an improved presentation of this paper. The first author was partially supported by NSFC 11871272. The second author was partially supported by NSFC 12071227 and 12371369, and the National Key Research and Development Program of China 2020YFA0713803. The third author was partially supported by NSFC 12171349 and 12271389.